Subject: Minutes, APO User's Committee, 12/13/99

From: strauss@astro.Princeton.EDU

Submitted: Tue, 14 Dec 1999 12:03:29 -0500 (EST)

Message number: 399 (previous: 398, next: 400 up: Index)

  Apache Point Observatory 3.5m User's Committee Meeting
		December 13, 1999

Attending: Jeff Brown, Lew Hobbs, Jon Holtzman, Bruce Gillespie, Ed
Turner, Michael Strauss

Not attending: Alan Uomoto, Chris Stubbs. 

Agenda: Status of the shutdown
	Issues from the report to the Board of Governors

***********************Status of telescope******************************
Turner, Gillespie: As of Friday, things were being reassembled; ready
to go on the sky for engineering Saturday or Sunday night, and going
into Science operations on the 17th.  These delays, to improve the
secondary support system in light of the problems with the Sloan
secondaries, have led to an intense engineering effort.
Unfortunately, the use of piezo-electrics as pressure sensors (to feed
limit switches), has failed for a whole variety of reasons.  There was
also hope to use piezos for the fast tip-tilt; it remains unclear
whether or not this will work.

  But mechanical hardstops, and galil box controllers, have all been
put in the redesigned secondary support system, which should make
things quite a bit safer than they were before. 

  As of Monday, December 13, the telescope is back together.
Unfortunately, one of the encoders on the tertiary is not functioning;
they are trying to figure this out now.  Given that the telescope has
been down for four months, it is likely that we'll find a number of
such minor glitches as we try to get back on the air.  Luckily, the
weather for tonight, and indeed, the rest of the week, is likely to be
good. [EDITOR'S NOTE: The encoder problem has now been fixed; the
telescope actually did do engineering tests on the sky last night
(i.e., the night of Monday, December 13).]

What are we doing for December 31, Y2K problems and all that?  Note
that in unusual circumstances (e.g., power outages, network going down
due to the millenium bug), the observatory is happy to do service
observing (i.e., by phone, with finding charts, written instructions, etc.).

  Turner will post a note to apo35-general describing the shutdown
history: why it has been so long, and why it has been extended so many
times.  Briefly, the cracks in the primary were discovered just as we
were coming out of the usual August shutdown; once we were on top of
this, the problems with the SDSS secondary came to light, which made
us seriously concerned for the safety of our secondary.  The risk to
the secondary is much diminished now, but not completely negligible.
There has been great progress in getting a final secondary support
design for the SDSS; we hope to use these ideas in the new secondary
top end, to be built over the summer. 

It would have been great if we could have used this downtime to do
much-needed work on the instruments (especially the DIS upgrade).
Part of the problem is that many of the same people who would work on
the instruments were working full-time on the telescope itself.  Also,
we didn't know at the beginning how long the shutdown would be, making
it difficult to make plans for work on the instruments. 

It is worth saying that the delays in getting the telescope back on
line are *not* due to people being distracted by the Sloan problems;
the people paid to work on the 3.5m have indeed been working full-time
on the 3.5m.

Under normal circumstances, in working on any major engineering task
like that we've been tackling, we would first do all relevant
preparatory work while the telescope was still running.
Unfortunately, given the emergency faced here, this was clearly not
possible, which is one reason the telescope has been down for so

  One piece of good news: the telescope with the new secondary did
observe the sky on a single night in late October.  Jeff Morgan has
analyzed the Shack-Hartman data taken; it implies that the telescope
should be able to deliver 0.5" images.  Although the exact calibration
of this number is a bit uncertain, it is clear that it is several
tenths of an arcsec better than it was with the old secondary.

Presentation to the ARC Board of Governors

Viewgraphs may be found at website:

Interesting plots/viewgraphs:
  -Number of APO publications as a function of time; a rapid increase recently!
  -DIMM measurements of local seeing for a typical night. 
  -Weather statistics, individual institution usage statistics.
   This leads to a discussion of observing efficiency, and how one
   measures this.  
  -Detailed diagrams about the cracks in the primary mirror. 
  -Was the image image quality poorer in 1999 than 1998?  Might this
   be due primarily to poor weather?  Known problems with the primary
   support system (which have since been fixed)?  DIMM should allow us
   to quantify this better.  
  -Shack-Hartmann diagrams for the new and old secondary. 
  -Hours of usage of each instrument. 
  -Summary of status of 3-year plan. 
  -Lists of published papers.  Note the official acknowledgements that
  should be included in any 3.5m paper:
  "Based on observations obtained with the Apache Point Observatory
3.5-meter telescope, which is owned and operated by the Astrophysical
Research Consortium" 

  The BoG approved the budget as requested. In particular, they
approved funding the next round of engineering tasks/improvement plans
for the observatory.  They did not approve funding for the aluminizing
facility, but encouraged us to explore creative financing arrangements
for it. 

  There is some concern that site-wide issues (telephone systems,
observatory vehicles), shared between the 3.5m and the SDSS, are
falling between the cracks. 

  Note that we don't have any major shutdowns planned over the next
two quarters. 

  Agenda items for next meeting from Jon Holtzman: 
   Status of the DIS upgrade: how are we going to make it move ahead?
Will it include an optics upgrade?  The UV throughput upgrade?  A new
camera for the DIS slitviewer? 
   Discussion of improving observing efficiency. 

  Previous month's minutes are approved. 

  The next meeting will be held on January 10 at 11:30 AM East Coast

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 399 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to